OpenStudy (anonymous):

2(x + 3) + -3 = 2x True False

OpenStudy (sasogeek):

2(x+3)-3=2x 2x+6-3=2x 2x+3=2x FALSE

OpenStudy (sasogeek):

\[2x+3\neq2x\]